### v2

# I-INR: Iterative Implicit Neural Representations

Abstract
Implicit Neural Representations (INRs) have revolutionized signal processing and computer vision by modeling signals as continuous, differentiable functions parameterized by neural networks. However, their inherent formulation as a regression problem makes them prone to regression to the mean, limiting their ability to capture fine details, retain high-frequency information, and handle noise effectively. To address these challenges, we propose Iterative Implicit Neural Representations (I-INRs) a novel plug-and-play framework that enhances signal reconstruction through an iterative refinement process. I-INRs effectively recover high-frequency details, improve robustness to noise, and achieve superior reconstruction quality. Our framework seamlessly integrates with existing INR architectures, delivering substantial performance gains across various tasks. Extensive experiments show that I-INRs outperform baseline methods, including WIRE, SIREN, and Gauss, in diverse computer vision applications such as image restoration, image denoising, and object occupancy prediction.
Figure 1: Effectiveness of the proposed methods across multiple tasks, including image fitting, Super Resolution (2$\times$ scale), and denoising, compared to the baseline representative INR method (SIREN [39]). Our novel Iterative-INR method (I-SIREN) consistently improves detail preservation, fidelity, and high-frequency reconstruction across all tasks, outperforming the baseline.
1 Introduction
Implicit Neural Representations (INRs) have emerged as a transformative approach in signal representation, shifting from traditional discrete grid-based methods to continuous coordinate-based models [39]. By leveraging neural networks, typically multi-layer perceptrons (MLPs), INRs map spatial or temporal coordinates directly to signal attributes such as pixel intensity, colour, or 3D occupancy. This continuous representation inherently offers resolution independence, compact encoding, and seamless interpolation, making INRs highly versatile across a wide range of domains including computer vision [5, 35, 39, 24, 18], and beyond [42, 37, 49, 31]. Initially introduced for tasks like 3D shape reconstruction [24, 26, 41] and novel view synthesis [17], INRs have since been applied to diverse applications including image fitting [23], super-resolution [9, 34], and solving inverse problems [3, 44, 5].
Despite advancements, INRs continue to face substantial challenges in capturing high-frequency details, maintaining robustness against noise, and handling incomplete data [36]. Typically optimized with L1 or L2 loss [39], INRs exhibit a bias toward low-frequency components due to the regression to the mean effect, leading to the loss of high-frequency details and reduced noise robustness [9]. Recent works have tried to solve this problem through positional encoding techniques  [43] or by incorporating new activation functions  [36, 39, 23]. Positional encoding techniques [43] incorporate various orthogonal Fourier [11], while periodic activations, such as sine functions, allow MLPs to better model high-frequency signals compared to conventional ReLU activations [39]. However, these methods are often based on the assumption of clean, noise-free data, which limits their applicability in real-world scenarios where noise and missing information are prevalent [36]. Advanced activation functions like complex Gabor wavelets have been proposed to enhance noise robustness [36], but these approaches still fall short when handling the complexities of challenging, real-world datasets. Despite these efforts, there remains significant potential for further improving high-frequency detail preservation, representational accuracy, and reconstruction quality in INRs.
To address these limitations, drawing inspiration from iterative models [9, 33, 15], we introduce Iterative Implicit Neural Representations (I-INRs), a plug-and-play framework that integrates seamlessly with existing implicit architectures. Unlike traditional single-shot INRs, I-INRs reconstruct signals progressively over multiple iterations, refining details, enhancing reconstruction quality, and improving noise robustness as shown in Figure 1. Our experiments demonstrate that I-INRs deliver superior reconstructions with enhanced detail, generalization, and minimal computational overhead, outperforming single-shot methods like WIRE [36], SIREN [39], and Gauss [30] across tasks such as image fitting, Super-Resolution (SR), denoising, and 3D occupancy reconstruction.
Our main contributions are.
  • We present Iterative Implicit Neural Representations, called I-INRs, a novel framework for INRs that reconstructs signals iteratively, effectively capturing high-frequency details, enhancing robustness to noise, and improving generalization.
  • I-INRs are designed as a plug-and-play framework, compatible with existing implicit architectures, enabling seamless integration and adoption in various applications.
  • We validate the effectiveness of I-INRs through extensive experiments across multiple tasks, demonstrating superior performance over traditional single-shot INRs in terms of detail reconstruction and robustness.
2 Related Work
INRs [39] map coordinates to signal values using neural networks, eliminating the need for discretization and thus offering flexibility in applications such as image, video processing [14, 5, 46, 4, 7], and 3D rendering [24, 32, 13]. However, INRs often suffer from spectral bias [29, 43], struggling to capture high-fidelity details of complex signals. This limitation affects applications requiring fine-grained detail, such as image SR, 3D reconstruction, and scene modeling. To address this challenge, various approaches have been proposed, including spatial encoding with frequency-based representations, polynomial decomposition [28, 38], and high-pass filtering [11], all of which emphasize high-frequency components. Beyond encoding techniques, the choice of activation functions also plays a crucial role in mitigating spectral bias. Standard activations like ReLU fail to capture high-frequency components due to their limited representational capacity [39]. To overcome this, alternative activation functions, such as sinusoidal [39] and Gaussian [30] activations, have been proposed. Additionally, approaches like WIRE [36] and FINER [23] introduce advanced activation strategies to enhance spectral bias tuning, with WIRE [36] leveraging Gabor wavelets and FINER [23] utilizing variable-periodic functions. INCODE [19] uses a task-specific pre-trained model that dynamically adjusts key parameters of activation function to enhance reconstruction performance. DINER [47] utilizes a hash map to unevenly map input coordinates onto a feature vector, optimizing the spatial frequency distribution and enabling faster, more precise reconstructions. While advanced encoding strategies and activation functions help mitigate spectral bias, they may not fully address the challenges of wide-bandwidth signals and large-scale data  [21, 35]. To scale INRs for wide-bandwidth signals, hierarchical and multi-scale designs like Progressive Implicit Networks (PINs) [21] and MINER [35] leverage hierarchical encoding and Laplacian pyramids, enabling efficient multi-scale detail capture while reducing computational overhead.
In contrast with single shot models, iterative models, such as diffusion models [15, 33, 9], generate high-quality samples by inverting known degradation processes. These models are widely used for diverse image processing tasks, including image restoration and translation [34, 33, 9, 16, 8, 6]. Furthermore, diffusion models have also been applied to 3D reconstruction tasks from text prompts or 2D images [27, 22]. Inspired by the strong detail reconstruction capabilities of iterative models, we propose a novel INR framework that reconstructs signals implicitly over multiple steps. Our method is a flexible, plug-and-play approach that can be integrated into any existing INR framework. However, for the scope of this paper, we specifically focus on WIRE [36], Gauss [30], and SIREN [39].
Figure 2: (a) The proposed architecture of the I-INR model. The framework consists of a Base INR, which can be any baseline architecture, denoted as Backbone, such as SIREN [39], WIRE [36], or Gauss [30]. Additionally, an Auxiliary Network incorporates feedback to refine representations, while a projection layer integrates features. The final output is obtained by combining the outputs of the Base INR and the projection layer, enhancing expressivity and reconstruction quality. (b) Iterative reconstruction processes of the proposed I-INR framework.
3 Proposed Method
In this section, we first provide an overview of INRs, followed by a detailed explanation of our proposed method, I-INR.
3.1 Preliminary: Standard INR
Let $g:X\subseteq\mathbb{R}^{p}\rightarrow Y\subseteq\mathbb{R}^{q}$ be a target function mapping $p$-dimensional inputs (such as pixel positions in a 2D image or volumetric domains in 3D scenes) to $q$-dimensional outputs (such as color or density values), so that $y=g(x)$ for ${}{x\in X}$. An INR [39] is a learnable neural function $f_{\theta}:X\rightarrow Y$, parameterized by $\theta$, aiming to approximate $\mathcal{I}(x)$ across $X$ [36]. The objective is to find $\theta$ such that:
$f_{\theta}(x)\approx\mathcal{I}(x),$ (1)
where $\mathcal{I}(x)$ represents an image function with pixel coordinates as $x$. A typical INR is trained by reducing the signal error in some $L_{p}$ metric,
$\min_{\theta}\mathbb{E}_{x,\mathcal{I}(x)}\|f_{\theta}(x)-\mathcal{I}(x)\|_{p}\approx\min_{\theta}\sum_{i}\|f_{\theta}(x^{i})-\mathcal{I}(x^{i})\|_{p},$ (2)
where $p=2$, the optimal minimum mean squared error (MMSE) solution is the conditional expectation as in Eq. (3)
[…]
This expectation leads to a signal that represents the average of all plausible reconstructions, often resulting in the loss of fine details. Such an approach limits the ability to capture high-frequency components and makes the reconstruction susceptible to noise. To overcome these limitations, we propose an iterative reconstruction approach that refines the signal progressively over multiple steps as formulated in Section 3.2. This iterative method enables a more precise reconstruction, enhancing frequency details and improving robustness to noise.
3.2 Iterative Implicit Neural Representations: I-INRs
I-INR leverages InDI [9] for its iterative setup. It reconstructs the signal in multiple steps, improving the reconstruction quality at each step. The function $g(x)_{t}$ at any time step $t$ can be defined as:
$g(x)_{t}=\mathcal{I}(x)(1-t)+\mathcal{Z}\,t,\quad t\in[0,1],$ (4)
where $\mathcal{Z}$ is sampled from a known distribution with the same dimensions as $\mathcal{I}(x)$. The process in Eq. (4) linearly interpolates between $\mathcal{I}(x)$ and $\mathcal{Z}$ using the time step $t$. At $t=0$, $g(x)_{0}$ equates to $\mathcal{I}(x)$, while at $t=1$, $g(x)_{1}$ is $\mathcal{Z}$. Thus, $g(x)_{t}$ smoothly transitions from $\mathcal{I}(x)$ to $\mathcal{Z}$ as $t$ moves from 0 to 1, providing a continuous interpolation between the two known states as shown in Figure 2(b).
Figure 3: Image fitting results for different non-linearities and their iterative extensions (prefix ”I”). Each column presents the reconstructed image (top) and residual map (bottom), highlighting reconstruction quality. PSNR values are reported, with iterative improvements in red. The rightmost column shows the ground truth.
During the iterative restoration process, we start from $\mathcal{Z}$ at $t=1$ and move toward the final reconstruction gradually. Let, $g(x)_{s}$ and $g(x)_{t}$ represent some intermediate states such that $s\leq t$. Then the conditional mean is given by:
$\mathbb{E}(g(x)_{s}\mid g(x)_{t})=\left(1-\frac{s}{t}\right)\mathbb{E}(g(x)_{0}\mid g(x)_{t})+\frac{s}{t}g(x)_{t}.$ (5)
When, $s=t-\delta$, then the estimated state $\hat{g}(x)_{t-\delta}$ is given by the Eq. 6 as:
$\begin{split}\hat{g}(x)_{t-\delta}&=\mathbb{E}(g(x)_{t-\delta}\mid g(x)_{t})\\ &=\frac{\delta}{t}\mathbb{E}(g(x)_{0}\mid\hat{g}(x)_{t})+\left(1-\frac{\delta}{t}\right)\hat{g}(x)_{t}.\end{split}$ (6)
Following Eq. (6), the reconstruction process begins at $g(x)_{1}=\mathcal{Z}$, with $t=1$. The final reconstruction is approached iteratively, taking steps of size $\delta$, by estimating the signal at each intermediate step.
Furthermore, taking the limit $\delta\to 0$ in Eq. (6) leads to the following ordinary differential equation (ODE)  [40, 9]:
$\frac{dg(x)_{t}}{dt}$ $=\lim_{\delta\to 0}\frac{g(x)_{t}-g(x)_{t-\delta}}{\delta}$ (7)
$=\frac{g(x)_{t}-\mathbb{E}(g(x)_{0}\mid g(x)_{t})}{t}.$
Here, the term $g(x)_{t}-\mathbb{E}[g(x)_{0}\mid g(x)_{t}]$ represents the residual error correction in the reconstruction process, where $\frac{1}{t}$ scales the update, allowing larger corrections in the early stages when $t$ is large, while finer adjustments take place as $t\to 0$, ensuring a smooth and gradual refinement toward $g(x)_{0}$ effectively preventing over-smoothing.
I-INR Training: Our proposed method aims to generate a progressively refined version of the image at each time step. To achieve this, we estimate a reconstructed version of the signal at each intermediate step $t$, specifically $\mathbb{E}(g(x)_{0}\mid g(x)_{t})$. This is achieved by training our neural function to predict the signal itself at any intermediate time step.
Given ${x},{I(x)}\sim p({x},{I(x)})$, $\mathcal{Z}\sim p(\mathcal{Z})$, and $t\sim p(t)$, the ideal objective function for this process is defined as:
$\min_{\theta}\mathbb{E}_{({x},{I(x)}),{\mathcal{Z}}}\mathbb{E}_{t}\|f_{\theta}({g(x)}_{t},{x},t)-{I(x)}\|_{2},$ (8)
where,
${g}(x)_{t}=\mathcal{I}(x)(1-t)+\mathcal{Z}t.$
To follow the regularity requirement mentioned in [9] for the iterative procedure, we add small noise to the intermediate latent state. The model with noise perturbation is given by:
$\tilde{g}_{t}(x)=I(x)(1-t)+\mathcal{Z}t+\epsilon\cdot n\cdot t,$ (9)
where, $\epsilon$ denotes a small constant that ensures a controlled amount of noise, and $n$ is a random variable sampled from a standard normal distribution $\mathcal{N}(0,1)$. Hence, our training objective becomes,
$\min_{\theta}\mathbb{E}_{({x},{I(x)}),{\mathcal{Z}}}\mathbb{E}_{t}\mathbb{E}_{n}\|f_{\theta}(\tilde{g}(x)_{t},{x},t)-{I(x)}\|_{2}.$ (10)
The detailed training procedure is provided in Algo. 1.
I-INR Reconstruction: During reconstruction, we start from an initial state $\mathcal{Z}$ and iteratively move towards the final reconstruction over $steps=\frac{1}{\delta}$. Typically, the known function $\mathcal{Z}$ is a simple initial state. Thus, we begin from $\mathcal{Z}$ and iteratively enhance it over time. The iterative implicit reconstruction process is outlined in Algo. 2.
Algorithm 1 I-INRs Model Training $f_{\theta}$
$\epsilon$
$(x,I(x))\sim p(x,I(x))$
$\mathcal{Z}\sim p(\mathcal{Z})$
repeat
     $t\sim\mathcal{U}(0,1)$
     $n\sim\mathcal{N}(0,I)$
     $\tilde{g}(x)_{t}=I(x)(1-t)+\mathcal{Z}t+\epsilon\cdot n\cdot t$
     Take a gradient descent step on
$\nabla_{\theta}\|f_{\theta}(\tilde{g}(x)_{t},x,t)-I(x)\|_{2}$
until converged
Algorithm 2 I-INRs Model Reconstruction
$\mathcal{Z}$, $f_{\theta}(\cdot,t)$, $\delta$, $x$
$\hat{g}(x)_{1}=\mathcal{Z}$
for $t\in\{1,1-\delta,1-2\delta,\dots,0\}$ do $\triangleright$ $t$ decreases from $1$ to 00 in steps of $\delta$
     $\hat{g}(x)_{t-\delta}=\frac{\delta}{t}f_{\theta}(\hat{g}(x)_{t},{x},t)+\left(1-\frac{\delta}{t}\right)\hat{g}(x)_{t}$
end for
return $\tilde{g}(x)_{0}$
Figure 4: Visual quality comparison of super-resolution results at 2× scale using various methods. The ground truth is compared against baseline INR methods SIREN [39], WIRE [36], Gauss [30], and their iterative counterparts. The iterative approaches consistently achieve sharper reconstructions with fewer artifacts, effectively preserving finer details and high-frequency structures.
Network Architecture: Figure 2(a) illustrates the I-INR architecture that integrates an Auxiliary Network and a Backbone (such as SIREN) to compute the output $I(x)_{t}$. The input $x$ is processed independently by both the Auxiliary Network and the Backbone, while the Auxiliary Network also incorporates feedback from the previous step. The output of the Auxiliary Network is concatenated with the features extracted by the Backbone, forming a joint representation. This concatenated feature then passes through a projection layer, which further refines the combined features. The final output is obtained as the dot product of the projection layer and the Backbone.
4 Experiments
To evaluate the effectiveness and robustness of our proposed method, we conduct extensive experiments across diverse tasks and backbones. Our evaluation encompasses regression tasks, such as 2D and 3D signal fitting, as well as generalization and robustness tasks, including image SR and image denoising. To demonstrate its comparative advantage, we benchmark our method against established baseline approaches, including SIREN [39], Gauss [30], and WIRE [36].
Implementation Settings: The proposed I-INR framework is implemented in PyTorch [25] and optimized using the Adam optimizer [10]. For all experiments, the noise parameter $\epsilon$ is empirically set to 0.1, and the inference $steps$ are fixed to 2 unless stated otherwise. The initial state $\mathcal{Z}$ is sampled from standard normal Gaussian distribution across all experiments. All baseline methods, including SIREN [39], WIRE [36], and Gauss [30], as well as their iterative counterparts, are initialized following the configurations specified in their respective papers. In I-INR, both the Auxiliary Network and the Projection layer employ lightweight MLPs to ensure computational efficiency while preserving expressivity.
Evaluation Metrics: To comprehensively evaluate our proposed method, we conduct extensive experiments using both distortion-based metrics, such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) [45], as well as perceptual metrics like Learned Perceptual Image Patch Similarity (LPIPS) [48].
4.1 Results
Our results consistently demonstrate that I-INR outperforms existing methods across diverse tasks, including image fitting, SR, and 3D occupancy. It achieves superior reconstruction quality and enhanced robustness compared to baseline approaches (SIREN, WIRE, Gauss).
Figure 5: Visual comparison of denoising results using various methods. The ground truth is compared against noisy input, baseline methods SIREN [39], WIRE [36], and Gauss [30], as well as their iterative counterparts. I-INR demonstrates superior artifact reduction and detail preservation compared to their non-iterative counterparts.
Table 1: Image fitting results on the Kodak dataset comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts. The iterative approach outperforms baseline single-shot architectures by achieving higher reconstruction quality and enhanced fidelity.
SIREN [39] WIRE [36] Gauss [30]
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Baseline 34.57 0.931 32.15 0.898 31.33 0.880
Ours 37.53 0.961 33.73 0.924 31.93 0.884
Image Fitting: For image fitting, we employ SIREN [39], WIRE [36], and Gauss [30], along with their iterative variants. Each network utilizes a three-layer MLP with 300 neurons per layer. In the case of iterative models, this corresponds to the Backbone INR illustrated in Figure 2(a). The networks are trained on the full-resolution Kodak dataset (768$\times$512 or 512$\times$768) [20], and the results, averaged across the entire dataset, are presented in Table 1. Our findings demonstrate that the proposed iterative models significantly outperform their one-shot counterparts. Figure 3 provides a qualitative comparison of all networks and their iterative adaptations on the kodim19 image. The error maps of baseline models exhibit homogeneous noise across the image, with more pronounced errors in high-frequency regions. In contrast, the iterative approaches yield smoother reconstructions with reduced errors throughout the image, including high-frequency areas. This ultimately results in a substantial improvement in fidelity, with PSNR gains of up to 2.7 dB observed for I-SIREN compared to its one-shot equivalent.
Image Super-resolution: We evaluate our I-INR framework against traditional one-shot INRs on SR tasks using 40 images from the DIV2K dataset [1]. The model is trained exclusively on 2×2$\times$2 × SR with a two-layer MLP of 256 neurons per layer. To assess generalization, we evaluate I-INR on both 2×2$\times$2 × and 4×4$\times$4 × SR, despite training only on 2×2$\times$2 ×. Table 2 compares its performance with single-shot baselines, demonstrating that our approach consistently enhances fidelity (measured as PSNR) and perceptual quality (measured as LPIPS) across all architectures. Figure 4 visualizes the butterfly image for 2×2$\times$2 × SR (please refer to supplementary material for visualization at 4×4$\times$4 × scale) , comparing each baseline with its iterative counterpart. The proposed approach consistently outperforms baseline methods, generating sharper reconstructions with fewer artifacts across all non-linearities. This improvement stems from the iterative process, which progressively refines the reconstruction, enabling more precise recovery of high-frequency details and allowing the model to generalize effectively beyond its training scale.
Table 2: Super-resolution (SR) results for 2×2$\times$2 × and 4×4$\times$4 × evaluations using a model trained exclusively on 2×2$\times$2 × SR. The table compares SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts. PSNR (higher is better) and LPIPS (lower is better) are reported, with the best result in each column highlighted in bold.
Scale Method SIREN [39] WIRE [36] Gauss [30]
Factor PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
2$\times$ Baseline 26.77 0.414 26.14 0.457 25.19 0.538
Ours 27.64 0.367 27.21 0.388 26.82 0.363
4$\times$ Baseline 25.03 0.597 24.57 0.6179 23.85 0.673
Ours 25.53 0.575 25.78 0.496 25.18 0.620
Table 3: Image denoising results on 40 images sampled from the DIV2K dataset [1], comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts. The best values in each column are highlighted in bold.
SIREN [39] WIRE [36] Gauss [30]
PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
Baseline 23.86 0.604 23.32 0.746 23.10 0.783
Ours 25.59 0.540 24.76 0.490 24.20 0.533
Figure 6: Visualization of 3D occupancy reconstruction results using various methods and their iterative counterparts. The ground truth is compared to baseline models (SIREN [39], WIRE [36], Gauss [30]) and their iterative counterparts. Iterative methods demonstrate superior reconstruction quality, achieving higher Intersection over Union (IOU) scores and capturing finer geometric details.
Image Denoising: To assess the robustness of I-INRs in modeling noisy signals, we utilize 40 high-resolution color images from the DIV2K dataset [1] and introduce Poisson-distributed noise to each pixel, with a maximum mean photon count of 30 and a readout noise level of 2, following  [36]. The same architecture and training procedure as SR is employed. Table 3 presents quantitative results comparing the denoising performance of SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts. Across all three methods, the iterative models consistently achieve significant improvements in both PSNR and LPIPS over their single-shot baselines. Figure 5 provides a qualitative comparison of denoising performance for SIREN [39], WIRE [36], and Gauss [30] alongside their iterative versions. The iterative approach significantly enhances reconstruction quality across all architectures, achieving up to +3.25 dB PSNR improvement (for SIREN).
Refer to caption Refer to caption
(a) (b)
Figure 7: On the Kodak dataset for image fitting: (a) PSNR vs. training FLOPs comparison between SIREN [39] and I-SIREN. (b) PSNR vs. training iterations for I-SIREN with different learning rates.
Figure 8: Analysis of I-SIREN’s performance for super-resolution at 2$\times$ resolution across different iteration $steps$. (a) I-SIREN achieves peak PSNR at $steps=4$, where further increasing time steps enhances perceptual quality but reduces fidelity due to the fidelity-perception tradeoff. (b) Increasing the number of steps consistently improves perceptual quality. (c) PSNR reaches its highest performance at $steps=4$, and while additional iterations slightly reduce PSNR, it remains superior to the baseline.
3D Occupancy: The 3D occupancy task aims to reconstruct 3D shapes by modeling the occupancy of points in space. We conduct experiments on the Armadillo, Dragon, and Thai Statue 3D scenes [12], utilizing a two-layer MLP with 256 neurons per layer for both the baseline INRs and their iterative counterparts. Table 4 presents the averaged Intersection over Union (IoU) results, while Figure 6 visualizes the reconstructions achieved using different nonlinearities and their iterative variants. Notably, I-INRs outperform their single-shot baselines, generating sharper and more precise reconstructions. The iterative approach enhances detail preservation, particularly in complex regions. For instance, I-SIREN reconstructs significantly finer details at the base of the statue compared to the baseline SIREN, demonstrating its ability to model intricate spatial structures more effectively.
Table 4: Averaged Intersection over Union (IoU) results comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts on the Armadillo, Dragon, and Thai Statue 3D scenes.
SIREN[39$\uparrow$ WIRE[36$\uparrow$ Gauss[30$\uparrow$
Baseline 0.9840 0.9917 0.9855
Ours 0.9934 0.9950 0.9967
4.2 Ablation Studies
Initial State: I-INR reconstructs a signal iteratively over multiple steps, starting from an initial state $\mathcal{Z}$. The choice of this initial state can significantly impact the reconstruction process. To assess its effect, we conduct extensive experiments with different types of initial states. Empirically, $\mathcal{Z}$ sampled from a standard normal Gaussian distribution $\mathcal{N}(0,1)$ consistently outperforms all-zeros and all-ones initial states across different non-linearities for the image fitting task. The results, summarized in Table 5, highlight the superiority of noise-based initial states in achieving higher PSNR and SSIM values.
Table 5: Impact of different initial states ($\mathcal{Z}$) on image fitting performance using the I-INR framework on the Kodak dataset. Noise-based initial states consistently achieve the best results across various non-linearities.
I-SIREN I-WIRE I-Gauss
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Noise 37.53 0.961 33.73 0.924 31.93 0.884
Ones 37.42 0.961 33.33 0.921 30.96 0.867
Zeros 37.47 0.961 33.39 0.921 31.22 0.882
Training and Inference Complexity: Our proposed I-INR method achieves faster convergence than the non-iterative baseline while maintaining similar training complexity, as shown in Figure 7a. The results are averaged over the Kodak dataset for the image fitting task. Figure 7b further illustrates I-SIREN’s training performance under different learning rates.
While I-SIREN converges faster, this improvement comes at a slightly higher per-iteration computational cost, requiring 107.2 GFLOPS compared to 106.8 GFLOPS for SIREN. Additionally, since I-SIREN reconstructs images iteratively, the total inference cost scales with the number of inference steps. In our experiments, we consistently use $steps=2$, resulting in a total inference complexity of 214.4 GFLOPS for I-SIREN, compared to 106.8 GFLOPS for SIREN.
Number of Layers: For the image fitting task on the Kodak dataset, we investigate the impact of increasing the number of layers in SIREN and its iterative counterpart, I-SIREN. The results, presented in Table 6, show that while the performance of SIREN saturates beyond a certain depth, specifically at five layers, I-SIREN continues to improve as more layers are added. Notably, a three-layer I-SIREN outperforms a five-layer SIREN, demonstrating the efficiency of the iterative approach.
Table 6: Impact of varying layer depths on the image fitting task using SIREN [39] and I-SIREN on the Kodak dataset. I-SIREN consistently outperforms SIREN [39], with a three-layer I-SIREN achieving higher PSNR than a five-layer SIREN [39].
Layers I-SIREN SIREN
PSNR Params (k) PSNR Params (k)
3 37.53 274 34.62 273
4 38.71 364 36.60 363
5 39.01 455 36.49 453
Analysis of Reconstruction Steps: We analyze the impact of reconstruction steps on 2$\times$ super-resolution using 40 images sampled from the DIV2K dataset, as shown in Figures 8b and 8c. Our findings reveal that the model achieves its highest PSNR at $steps=4$, with the most significant gain occurring when $steps$ increases from 1 to 2. Beyond $steps=2$, the PSNR begins to plateau. Notably, although PSNR decreases for $steps>4$, the perceptual quality (measured by LPIPS) continues to improve. This behavior aligns with the Perception-Distortion Tradeoff [2], where introducing additional high-frequency details enhances realism but slightly deviates from the ground-truth, thus lowering distortion-based metrics such as PSNR. Figure 8a presents a visual comparison of SIREN and I-SIREN on an image from DIV2K dataset at various time steps. I-SIREN reconstructions become progressively more vibrant and perceptually appealing as step increases, confirming that additional refinement steps are beneficial for capturing finer textures and high-frequency components. This trend holds for other reconstruction tasks as well, as detailed in our supplementary material.
5 Conclusion
This work introduces a novel I-INR framework that progressively reconstructs a signal. The proposed method serves as a plug-and-play solution compatible with existing INR methods, significantly enhancing their performance. Through extensive experiments, we demonstrate that I-INR effectively captures high-frequency details, improves noise robustness, and achieves superior reconstruction quality across various tasks and nonlinearities.
Supplementary Material
6 Additional Results
6.1 Additional Quantative Results
Table 7 presents a detailed comparison of our proposed method against the baseline across image fitting, denoising, and SR at 2×2$\times$2 × for all three evaluation metrics (PSNR, SSIM, and LPIPS). Our iterative approach consistently enhances performance across tasks and different baselines, achieving superior reconstruction quality.
6.2 Additional Qualitative Comparison
We also present an additional qualitative comparison of our proposed method against SIREN, Gauss, and WIRE. Figures 9101112, illustrates a comparative analysis for image fitting, SR at 2×2$\times$2 × and 4×4$\times$4 × scales, and image denoising, demonstrating the improvements achieved by our approach over baseline INRs. Our proposed method consistently outperforms single-shot baselines, demonstrating superior reconstruction quality across various models and tasks.
7 Additional Ablation Studies
Additional Reconstruction Steps Analysis: For SR at 2×2$\times$2 ×, we further analyze the reconstruction quality of I-SIREN over multiple steps, comparing it to SIREN and the ground truth. As shown in Figure 13, the results demonstrate that with an increasing number of steps, details become more pronounced, and more high-frequency information is incorporated into the reconstruction.
Additionally, We analyze the impact of reconstruction steps on image fitting using the Kodak dataset (see Figure 14). Our findings indicate that the model achieves its highest PSNR at $steps=4$, with the most significant gain occurring between $steps=1$ and $steps=2$. Beyond $steps=2$, PSNR begins to plateau. Notably, even though PSNR decreases for $steps>4$, perceptual quality continues to improve.
We also analyze the effect of reconstruction steps on image denoising performance using 40 images from the DIV2K dataset (see Figure 15). Our findings indicate that PSNR peaks at $steps=2$ before gradually declining as $steps$ increases. Meanwhile, LPIPS improves until $step=4$ and stabilizes around step 5. Beyond $steps=4$, both PSNR and LPIPS degrade as I-INR begins reconstructing noise present in the data. At higher iterations, the model inadvertently reconstructs these distortions, leading to a decline in performance.
Feature Fusion: We conduct ablation studies on the final feature-fusion mechanism for signal reconstruction, evaluating two strategies that combine the outputs of the Backbone and Projection modules (Figure 2a of main manuscript):
[…]
In Multiplicative fusion, the final output $\mathcal{I}(x)_{t}$ is obtained via element-wise multiplication of the Backbone and Projection outputs. The Adaptive method, conversely, uses the interpolation factor $t$ to smoothly shift emphasis from the Backbone output (dominant at $t=1$) toward the Projection output at higher values of $t$. Quantitative results comparing both fusion strategies for the image-fitting task are summarized in Table 8. Multiplicative fusion consistently outperforms Adaptive fusion, as evidenced by the results. Therefore, we adopt Multiplicative fusion as the standard approach for our method.
Table 7: Experimental results for Denoising, Image Fitting, and Super Resolution for Baseline model with Ours, evaluating their performance across three tasks. The best values in each column are highlighted in bold.
SIREN Gauss WIRE
[…]
Image Fitting Baseline 34.57 0.931 0.080 31.33 0.880 0.147 32.15 0.898 0.120
Ours 37.53 0.961 0.032 31.93 0.884 0.124 33.73 0.924 0.077
Super Resolution Baseline 26.77 0.763 0.414 25.19 0.709 0.538 26.14 0.749 0.457
Ours 27.64 0.783 0.367 26.82 0.761 0.363 27.21 0.778 0.388
Denoising Baseline 23.86 0.658 0.604 23.09 0.604 0.783 23.32 0.596 0.746
Ours 25.59 0.619 0.540 24.20 0.612 0.533 24.76 0.645 0.490
Table 8: Comparison of fusion methods for the image-fitting task at $steps=2$ on the Kodak dataset with Gaussian initial state ($\mathcal{Z}\sim\mathcal{N}(0,1)$).
I-SIREN I-WIRE I-Gauss
[…]
Multiplicative 37.53 0.961 0.032 33.73 0.924 0.077 31.93 0.884 0.124
Adaptive 37.01 0.960 0.037 33.63 0.923 0.078 31.80 0.883 0.129
8 Training Setup
Table 9 presents the key hyperparameters for various experiments. These hyperparameters were chosen to maximize peak performance across different tasks.
Additionally, we used 2000 training iterations for image fitting, super-resolution (SR), and denoising tasks, while for object occupancy, we used 200 training iterations.
Table 9: Hyperparameter settings for I-INRs across different tasks.
Task Parameter I-Gauss I-SIREN I-WIRE
Image Fitting $\omega$ 52 7
$\sigma$ 18 - 13
Image SR $\omega$ 30 4
$\sigma$ 11 - 10
Image Denoising $\omega$ 55 10
$\sigma$ 18 - 16
3D Occupancy $\omega$ 55 10
$\sigma$ 17 - 20
Figure 9: Image fitting results for different non-linearities and their iterative extensions (prefix ”I”).
Figure 10: Visual quality comparison of super-resolution results at 2$\times$ scale using various methods. The ground truth is compared against baseline INR methods SIREN, WIRE, Gauss, and their iterative counterparts. The iterative approaches consistently achieve sharper reconstructions with fewer artifacts, effectively preserving finer details and high-frequency structures.
Figure 11: Visual quality comparison of super-resolution results at 4$\times$ scale using various methods. The ground truth is compared against baseline INR methods SIREN, WIRE, Gauss, and their iterative counterparts. The iterative approaches consistently achieve sharper reconstructions with fewer artifacts, effectively preserving finer details and high-frequency structures.
Figure 12: Visual comparison of denoising results using various methods. The ground truth is compared against noisy input, baseline methods SIREN, WIRE, and Gauss, as well as their iterative counterparts. I-INR demonstrates superior artifact reduction and detail preservation compared to their non-iterative counterparts.
Figure 13: Analysis of I-SIREN’s performance for super-resolution at 2$\times$ resolution across different steps. I-SIREN achieves peak PSNR at $steps=4$, where further increasing steps enhances perceptual quality.
Refer to caption Refer to caption
(a) (b)
Figure 14: Analysis of I-SIREN’s performance for image fitting across different iteration steps. (a) Increasing the number of steps consistently improves perceptual quality. (b) PSNR reaches its highest performance at $steps=4$, and while additional iterations slightly reduce PSNR, it remains superior to the baseline.
Refer to caption Refer to caption
(a) (b)
Figure 15: Analysis of I-SIREN’s performance for denoising across different reconstruction steps. (a) LPIPS improves until step 4 and stabilizes around step 5. (b) PSNR reaches its peak at $steps=2$, but further iterations degrade performance.

### v3

@@
 Implicit Neural Representations (INRs) have revolutionized signal processing and computer vision by modeling signals as continuous, differentiable functions parameterized by neural networks.
-However, their inherent formulation as a regression problem makes them prone to regression to the mean, limiting their ability to capture fine details, retain high-frequency information, and handle noise effectively.
-To address these challenges, we propose Iterative Implicit Neural Representations (I-INRs) a novel plug-and-play framework that enhances signal reconstruction through an iterative refinement process.
-I-INRs effectively recover high-frequency details, improve robustness to noise, and achieve superior reconstruction quality.
-Our framework seamlessly integrates with existing INR architectures, delivering substantial performance gains across various tasks.
-Extensive experiments show that I-INRs outperform baseline methods, including WIRE, SIREN, and Gauss, in diverse computer vision applications such as image restoration, image denoising, and object occupancy prediction.
-Figure 1: Effectiveness of the proposed methods across multiple tasks, including image fitting, Super Resolution (2$\times$ scale), and denoising, compared to the baseline representative INR method (SIREN [39]).
-Our novel Iterative-INR method (I-SIREN) consistently improves detail preservation, fidelity, and high-frequency reconstruction across all tasks, outperforming the baseline.
+However, INRs are prone to the spectral bias problem, limiting their ability to retain high-frequency information, and often struggle with noise robustness.
+Motivated by recent trends in iterative refinement processes, we propose Iterative Implicit Neural Representations (I-INRs).
+This novel plug-and-play framework iteratively refines signal reconstructions to restore high-frequency details, improve noise robustness, and enhance generalization, ultimately delivering superior reconstruction quality.
+I-INRs integrate seamlessly into existing INR architectures with only a 0.5–2% increase in parameters.
+During reconstruction, the iterative refinement adds just 0.8–1.6% additional FLOPs over the baseline while delivering a substantial performance boost of up to +2.0 PSNR.
+Extensive experiments demonstrate that I-INRs consistently outperform WIRE, SIREN, and Gauss across various computer vision tasks, including image fitting, image denoising, and object occupancy prediction.
+The code is available at https://github.com/optimizer077/I-INR.
+[…]
 1 Introduction
-Implicit Neural Representations (INRs) have emerged as a transformative approach in signal representation, shifting from traditional discrete grid-based methods to continuous coordinate-based models [39].
+Implicit Neural Representations (INRs) have emerged as a transformative approach in signal representation, shifting from traditional discrete grid-based methods to continuous coordinate-based models (sitzmann2020implicit).
 By leveraging neural networks, typically multi-layer perceptrons (MLPs), INRs map spatial or temporal coordinates directly to signal attributes such as pixel intensity, colour, or 3D occupancy.
-This continuous representation inherently offers resolution independence, compact encoding, and seamless interpolation, making INRs highly versatile across a wide range of domains including computer vision [5, 35, 39, 24, 18], and beyond [42, 37, 49, 31].
-Initially introduced for tasks like 3D shape reconstruction [24, 26, 41] and novel view synthesis [17], INRs have since been applied to diverse applications including image fitting [23], super-resolution [9, 34], and solving inverse problems [3, 44, 5].
-Despite advancements, INRs continue to face substantial challenges in capturing high-frequency details, maintaining robustness against noise, and handling incomplete data [36].
-Typically optimized with L1 or L2 loss [39], INRs exhibit a bias toward low-frequency components due to the regression to the mean effect, leading to the loss of high-frequency details and reduced noise robustness [9].
-Recent works have tried to solve this problem through positional encoding techniques  [43] or by incorporating new activation functions  [36, 39, 23].
-Positional encoding techniques [43] incorporate various orthogonal Fourier [11], while periodic activations, such as sine functions, allow MLPs to better model high-frequency signals compared to conventional ReLU activations [39].
-However, these methods are often based on the assumption of clean, noise-free data, which limits their applicability in real-world scenarios where noise and missing information are prevalent [36].
-Advanced activation functions like complex Gabor wavelets have been proposed to enhance noise robustness [36], but these approaches still fall short when handling the complexities of challenging, real-world datasets.
-Despite these efforts, there remains significant potential for further improving high-frequency detail preservation, representational accuracy, and reconstruction quality in INRs.
-To address these limitations, drawing inspiration from iterative models [9, 33, 15], we introduce Iterative Implicit Neural Representations (I-INRs), a plug-and-play framework that integrates seamlessly with existing implicit architectures.
-Unlike traditional single-shot INRs, I-INRs reconstruct signals progressively over multiple iterations, refining details, enhancing reconstruction quality, and improving noise robustness as shown in Figure 1.
-Our experiments demonstrate that I-INRs deliver superior reconstructions with enhanced detail, generalization, and minimal computational overhead, outperforming single-shot methods like WIRE [36], SIREN [39], and Gauss [30] across tasks such as image fitting, Super-Resolution (SR), denoising, and 3D occupancy reconstruction.
-Our main contributions are.
+This continuous representation inherently offers resolution independence, compact encoding, and seamless interpolation, making INRs highly versatile across a wide range of domains, including computer vision (chen2021learning; saragadam2022miner; sitzmann2020implicit; mildenhall2021nerf; jiang2023coordinate), and beyond (sun2021coil; shen2021non; zhong2019reconstructing; reed2021implicit).
+Initially introduced for tasks like 3D shape reconstruction (mildenhall2021nerf; peng2020convolutional; srinivasan2021nerv) and novel view synthesis (Irshad_2023_ICCV), INRs have since been applied to diverse applications including image fitting (liu2023finer), super-resolution (delbracioinversion; saharia2022image), and solving inverse problems (cha2024descanning; wang2019spatial; chen2021learning).
+Despite advancements, INRs continue to face substantial challenges in capturing high-frequency details, maintaining robustness against noise, and handling incomplete data (saragadam2023wire).
+Typically optimized with L1 or L2 loss (sitzmann2020implicit), INRs inherently exhibit a spectral bias that favors low-frequency components, often at the expense of fine-grained details (rahaman2019spectral).
+To mitigate this, recent approaches have introduced positional encoding schemes (tancik2020fourier; fathony2020multiplicative) and alternative activation functions (sitzmann2020implicit; saragadam2023wire; liu2023finer; gao2024h; thennakoon2025bandrc) to better capture high-frequency content.
+Positional encodings inject high-frequency signals through orthogonal Fourier bases (tancik2020fourier), while periodic activations such as sine functions enable MLPs to model high-frequency structures more effectively compared to traditional ReLU-based networks (sitzmann2020implicit).
+Nevertheless, these methods often assume clean and fully observed inputs, limiting their effectiveness in practical scenarios characterized by noise and occlusion (saragadam2023wire).
+More recent efforts, such as the integration of complex Gabor wavelet activations, have aimed to improve robustness under such conditions (saragadam2023wire), but challenges persist when scaling to real-world, noisy data.
+Consequently, there remains significant potential to enhance the fidelity, robustness, and generalization of INRs, particularly in preserving high-frequency signals under adverse conditions.
+To address these limitations, and drawing inspiration from iterative models (delbracioinversion; rissanen2022generative; ho2020denoising), we introduce Iterative Implicit Neural Representations (I-INRs) a plug-and-play framework that integrates seamlessly with existing implicit architectures.
+Unlike traditional single-shot INRs, I-INRs reconstruct signals progressively over multiple iterations, refining details, enhancing reconstruction quality, and improving noise robustness, as illustrated in Figure 1.
+The proposed framework features a novel architecture that incorporates existing implicit networks as a BackboneNet without major modifications, while introducing two lightweight modules, i.e FeedbackNet and FuseNet as shown in Figure 2(a).
+These blocks add only 0.5–2% to the total parameter count of the baseline model and support the iterative reconstruction process.
+Owing to this design, the iterative I-INR process incurs only 0.8–2% additional FLOPs over the baseline INR for two-step reconstruction.
+Our experiments demonstrate that I-INRs deliver superior reconstructions with enhanced detail, better generalization, and minimal computational overhead, consistently outperforming single-shot methods such as WIRE (saragadam2023wire), SIREN (sitzmann2020implicit), and Gauss (ramasinghe2022beyond) across tasks including image fitting, super-resolution, denoising, and 3D occupancy reconstruction.
+Our main contributions are as follows:
   • We present Iterative Implicit Neural Representations, called I-INRs, a novel framework for INRs that reconstructs signals iteratively, effectively capturing high-frequency details, enhancing robustness to noise, and improving generalization.
@@
 2 Related Work
-INRs [39] map coordinates to signal values using neural networks, eliminating the need for discretization and thus offering flexibility in applications such as image, video processing [14, 5, 46, 4, 7], and 3D rendering [24, 32, 13].
-However, INRs often suffer from spectral bias [29, 43], struggling to capture high-fidelity details of complex signals.
-This limitation affects applications requiring fine-grained detail, such as image SR, 3D reconstruction, and scene modeling.
-To address this challenge, various approaches have been proposed, including spatial encoding with frequency-based representations, polynomial decomposition [28, 38], and high-pass filtering [11], all of which emphasize high-frequency components.
-Beyond encoding techniques, the choice of activation functions also plays a crucial role in mitigating spectral bias.
-Standard activations like ReLU fail to capture high-frequency components due to their limited representational capacity [39].
-To overcome this, alternative activation functions, such as sinusoidal [39] and Gaussian [30] activations, have been proposed.
-Additionally, approaches like WIRE [36] and FINER [23] introduce advanced activation strategies to enhance spectral bias tuning, with WIRE [36] leveraging Gabor wavelets and FINER [23] utilizing variable-periodic functions.
-INCODE [19] uses a task-specific pre-trained model that dynamically adjusts key parameters of activation function to enhance reconstruction performance.
-DINER [47] utilizes a hash map to unevenly map input coordinates onto a feature vector, optimizing the spatial frequency distribution and enabling faster, more precise reconstructions.
-While advanced encoding strategies and activation functions help mitigate spectral bias, they may not fully address the challenges of wide-bandwidth signals and large-scale data  [21, 35].
-To scale INRs for wide-bandwidth signals, hierarchical and multi-scale designs like Progressive Implicit Networks (PINs) [21] and MINER [35] leverage hierarchical encoding and Laplacian pyramids, enabling efficient multi-scale detail capture while reducing computational overhead.
-In contrast with single shot models, iterative models, such as diffusion models [15, 33, 9], generate high-quality samples by inverting known degradation processes.
-These models are widely used for diverse image processing tasks, including image restoration and translation [34, 33, 9, 16, 8, 6].
-Furthermore, diffusion models have also been applied to 3D reconstruction tasks from text prompts or 2D images [27, 22].
+Spatial Encoding Techniques.
+INRs suffer from spectral bias (rahaman2019spectral; tancik2020fourier), struggling to capture high-fidelity details of complex signals.
+This limitation affects applications requiring fine-grained detail, such as a single image super resolution (SR), 3D reconstruction, and scene modeling.
+To address this challenge, various approaches have been proposed, including spatial encoding with frequency-based representations, polynomial decomposition (raghavan2023neural; singh2023polynomial), and high-pass filtering (fathony2020multiplicative), all of which emphasize high-frequency components.
+DINER (xie2023diner) applies a hash map to unevenly map input coordinates to feature vectors, optimizing spatial frequency distribution for faster, more accurate reconstructions.
+Role of Activations.
+Beyond encoding, activation functions critically influence spectral bias.
+Standard choices like ReLU cannot capture high-frequency components due to limited representational capacity (sitzmann2020implicit).
+To address this, alternatives such as sinusoidal (sitzmann2020implicit) and Gaussian (ramasinghe2022beyond) activations have been proposed.
+Activations tuned for high accuracy can suffer from reduced noise robustness; WIRE  (saragadam2023wire) leverages Gabor wavelets to provide enhanced robustness to noise.
+However, there remains room for improvement in achieving both high-frequency fitting and noise robustness  (saragadam2023wire).
+Iterative Methods.
+Iterative models have been very successful in image restoration and translation (saharia2022image; rissanen2022generative; delbracioinversion; hui2024microdiffusion; chu2025highly; chen2024image), video generation (ji2025layerflow), language modeling (nie2025large), 3D generation (qian2023magic123; liu2023zero), and more.
+In contrast with single-shot models, iterative models such as diffusion models (ho2020denoising; rissanen2022generative; delbracioinversion) produce high-quality samples by reversing degradation processes over multiple steps.
+Despite their widespread adoption and effectiveness, iterative models remain underexplored in the context of implicit neural representations.
 Inspired by the strong detail reconstruction capabilities of iterative models, we propose a novel INR framework that reconstructs signals implicitly over multiple steps.
-Our method is a flexible, plug-and-play approach that can be integrated into any existing INR framework.
-However, for the scope of this paper, we specifically focus on WIRE [36], Gauss [30], and SIREN [39].
+Our approach provides improved details and enhanced noise robustness, offering a flexible, plug-and-play solution that can be integrated into any existing INR framework.
+However, for the scope of this paper, we specifically focus on WIRE (saragadam2023wire), Gauss (ramasinghe2022beyond), and SIREN (sitzmann2020implicit).
 Figure 2: (a) The proposed architecture of the I-INR model.
-The framework consists of a Base INR, which can be any baseline architecture, denoted as Backbone, such as SIREN [39], WIRE [36], or Gauss [30].
-Additionally, an Auxiliary Network incorporates feedback to refine representations, while a projection layer integrates features.
-The final output is obtained by combining the outputs of the Base INR and the projection layer, enhancing expressivity and reconstruction quality.
-(b) Iterative reconstruction processes of the proposed I-INR framework.
-3 Proposed Method
-In this section, we first provide an overview of INRs, followed by a detailed explanation of our proposed method, I-INR.
-3.1 Preliminary: Standard INR
-Let $g:X\subseteq\mathbb{R}^{p}\rightarrow Y\subseteq\mathbb{R}^{q}$ be a target function mapping $p$-dimensional inputs (such as pixel positions in a 2D image or volumetric domains in 3D scenes) to $q$-dimensional outputs (such as color or density values), so that $y=g(x)$ for ${}{x\in X}$.
-An INR [39] is a learnable neural function $f_{\theta}:X\rightarrow Y$, parameterized by $\theta$, aiming to approximate $\mathcal{I}(x)$ across $X$ [36].
+The framework consists of a Backbone, which can be any baseline INR architecture.
+Additionally, a Feedback Net incorporates feedback to refine representations, while FuseNet integrates features.
+The final output is obtained by combining the outputs of the Backbone and the FuseNet, enhancing expressivity and reconstruction quality.
+(b) Iterative reconstruction process of the proposed I-INR framework.
+3 Preliminary
+3.1 Standard INR
+Let $g:X\subseteq\mathbb{R}^{p}\rightarrow Y\subseteq\mathbb{R}^{q}$ be a target function mapping $p$-dimensional inputs (such as pixel positions in a 2D image or volumetric domains in 3D scenes) to $q$-dimensional outputs (such as color or density values), such that $y=g(x)$ for $x\in X$.
+An INR (sitzmann2020implicit) is a learnable neural function $f_{\theta}:X\rightarrow Y$, parameterized by $\theta$, aiming to approximate $\mathcal{I}(x)$ across $X$.
 The objective is to find $\theta$ such that:
 $f_{\theta}(x)\approx\mathcal{I}(x),$ (1)
-where $\mathcal{I}(x)$ represents an image function with pixel coordinates as $x$. A typical INR is trained by reducing the signal error in some $L_{p}$ metric,
-$\min_{\theta}\mathbb{E}_{x,\mathcal{I}(x)}\|f_{\theta}(x)-\mathcal{I}(x)\|_{p}\approx\min_{\theta}\sum_{i}\|f_{\theta}(x^{i})-\mathcal{I}(x^{i})\|_{p},$ (2)
-where $p=2$, the optimal minimum mean squared error (MMSE) solution is the conditional expectation as in Eq. (3)
-[…]
-This expectation leads to a signal that represents the average of all plausible reconstructions, often resulting in the loss of fine details.
-Such an approach limits the ability to capture high-frequency components and makes the reconstruction susceptible to noise.
-To overcome these limitations, we propose an iterative reconstruction approach that refines the signal progressively over multiple steps as formulated in Section 3.2.
-This iterative method enables a more precise reconstruction, enhancing frequency details and improving robustness to noise.
-3.2 Iterative Implicit Neural Representations: I-INRs
-I-INR leverages InDI [9] for its iterative setup.
-It reconstructs the signal in multiple steps, improving the reconstruction quality at each step.
-The function $g(x)_{t}$ at any time step $t$ can be defined as:
-$g(x)_{t}=\mathcal{I}(x)(1-t)+\mathcal{Z}\,t,\quad t\in[0,1],$ (4)
-where $\mathcal{Z}$ is sampled from a known distribution with the same dimensions as $\mathcal{I}(x)$.
-The process in Eq.
-(4) linearly interpolates between $\mathcal{I}(x)$ and $\mathcal{Z}$ using the time step $t$.
-At $t=0$, $g(x)_{0}$ equates to $\mathcal{I}(x)$, while at $t=1$, $g(x)_{1}$ is $\mathcal{Z}$.
-Thus, $g(x)_{t}$ smoothly transitions from $\mathcal{I}(x)$ to $\mathcal{Z}$ as $t$ moves from 0 to 1, providing a continuous interpolation between the two known states as shown in Figure 2(b).
+where $\mathcal{I}(x)$ represents a true image function evaluated at coordinates $x$. A typical INR is trained by reducing the signal error in an $L_{p}$ metric:
+$\min_{\theta}\mathbb{E}_{x}\left[\|f_{\theta}(x)-\mathcal{I}(x)\|_{p}\right]\approx\min_{\theta}\sum_{i}\|f_{\theta}(x^{i})-\mathcal{I}(x^{i})\|_{p},$ (2)
+where $p=2$ for standard INR training, corresponding to the mean squared error.
+3.2 Inversion by Direct Iteration (InDI)
+Inversion by Direct Iteration (InDI) (delbracioinversion) addresses ill-posed inverse problems, such as image restoration, by iteratively refining an estimate of the unknown signal $b$ from an observed measurement $a$.
+Instead of solving the full inverse problem in one step, InDI interpolates between $a$ and $b$ via $b_{t}=ta+(1-t)b$, with $t\in[0,1]$, and uses a neural network $F_{\theta}$ to approximate the conditional expectation $\mathbb{E}[b\mid b_{t}]$.
+The network is trained to minimize:
+$\min_{\theta}\,\mathbb{E}_{a,b,t,n}\left[\,\|F_{\theta}(b_{t}+\epsilon_{t}n,\,t)-b\|_{1}\,\right],$
+where $\epsilon_{t}n$ is Gaussian noise. At inference, the estimate is updated recursively using the learned conditional mean, gradually solving a sequence of easier inverse problems from $a$ to $b$.
+4 Proposed Method
+In this section, we present a detailed explanation of our proposed methods, including a comprehensive description of the iterative plug-and-play framework.
+4.1 Iterative Implicit Neural Representations: I-INR
 Figure 3: Image fitting results for different non-linearities and their iterative extensions (prefix ”I”).
@@
 The rightmost column shows the ground truth.
-During the iterative restoration process, we start from $\mathcal{Z}$ at $t=1$ and move toward the final reconstruction gradually.
-Let, $g(x)_{s}$ and $g(x)_{t}$ represent some intermediate states such that $s\leq t$.
-Then the conditional mean is given by:
-$\mathbb{E}(g(x)_{s}\mid g(x)_{t})=\left(1-\frac{s}{t}\right)\mathbb{E}(g(x)_{0}\mid g(x)_{t})+\frac{s}{t}g(x)_{t}.$ (5)
-When, $s=t-\delta$, then the estimated state $\hat{g}(x)_{t-\delta}$ is given by the Eq. 6 as:
-$\begin{split}\hat{g}(x)_{t-\delta}&=\mathbb{E}(g(x)_{t-\delta}\mid g(x)_{t})\\ &=\frac{\delta}{t}\mathbb{E}(g(x)_{0}\mid\hat{g}(x)_{t})+\left(1-\frac{\delta}{t}\right)\hat{g}(x)_{t}.\end{split}$ (6)
-Following Eq.
-(6), the reconstruction process begins at $g(x)_{1}=\mathcal{Z}$, with $t=1$.
-The final reconstruction is approached iteratively, taking steps of size $\delta$, by estimating the signal at each intermediate step.
-Furthermore, taking the limit $\delta\to 0$ in Eq. (6) leads to the following ordinary differential equation (ODE)  [40, 9]:
-$\frac{dg(x)_{t}}{dt}$ $=\lim_{\delta\to 0}\frac{g(x)_{t}-g(x)_{t-\delta}}{\delta}$ (7)
-$=\frac{g(x)_{t}-\mathbb{E}(g(x)_{0}\mid g(x)_{t})}{t}.$
-Here, the term $g(x)_{t}-\mathbb{E}[g(x)_{0}\mid g(x)_{t}]$ represents the residual error correction in the reconstruction process, where $\frac{1}{t}$ scales the update, allowing larger corrections in the early stages when $t$ is large, while finer adjustments take place as $t\to 0$, ensuring a smooth and gradual refinement toward $g(x)_{0}$ effectively preventing over-smoothing.
-I-INR Training: Our proposed method aims to generate a progressively refined version of the image at each time step.
-To achieve this, we estimate a reconstructed version of the signal at each intermediate step $t$, specifically $\mathbb{E}(g(x)_{0}\mid g(x)_{t})$.
-This is achieved by training our neural function to predict the signal itself at any intermediate time step.
-Given ${x},{I(x)}\sim p({x},{I(x)})$, $\mathcal{Z}\sim p(\mathcal{Z})$, and $t\sim p(t)$, the ideal objective function for this process is defined as:
-$\min_{\theta}\mathbb{E}_{({x},{I(x)}),{\mathcal{Z}}}\mathbb{E}_{t}\|f_{\theta}({g(x)}_{t},{x},t)-{I(x)}\|_{2},$ (8)
-where,
-${g}(x)_{t}=\mathcal{I}(x)(1-t)+\mathcal{Z}t.$
-To follow the regularity requirement mentioned in [9] for the iterative procedure, we add small noise to the intermediate latent state. The model with noise perturbation is given by:
-$\tilde{g}_{t}(x)=I(x)(1-t)+\mathcal{Z}t+\epsilon\cdot n\cdot t,$ (9)
-where, $\epsilon$ denotes a small constant that ensures a controlled amount of noise, and $n$ is a random variable sampled from a standard normal distribution $\mathcal{N}(0,1)$.
-Hence, our training objective becomes,
-$\min_{\theta}\mathbb{E}_{({x},{I(x)}),{\mathcal{Z}}}\mathbb{E}_{t}\mathbb{E}_{n}\|f_{\theta}(\tilde{g}(x)_{t},{x},t)-{I(x)}\|_{2}.$ (10)
-The detailed training procedure is provided in Algo. 1.
-I-INR Reconstruction: During reconstruction, we start from an initial state $\mathcal{Z}$ and iteratively move towards the final reconstruction over $steps=\frac{1}{\delta}$.
-Typically, the known function $\mathcal{Z}$ is a simple initial state.
-Thus, we begin from $\mathcal{Z}$ and iteratively enhance it over time.
-The iterative implicit reconstruction process is outlined in Algo.
-2.
-Algorithm 1 I-INRs Model Training $f_{\theta}$
-$\epsilon$
-$(x,I(x))\sim p(x,I(x))$
-$\mathcal{Z}\sim p(\mathcal{Z})$
-repeat
-     $t\sim\mathcal{U}(0,1)$
-     $n\sim\mathcal{N}(0,I)$
-     $\tilde{g}(x)_{t}=I(x)(1-t)+\mathcal{Z}t+\epsilon\cdot n\cdot t$
-     Take a gradient descent step on
-$\nabla_{\theta}\|f_{\theta}(\tilde{g}(x)_{t},x,t)-I(x)\|_{2}$
-until converged
+I-INR reconstructs the signal over multiple steps, progressively improving the quality at each iteration.
+As shown in Figure 2(b), the process starts from an initial state $g(x)_{1}=\mathcal{Z}$ at $t=1$ and gradually approaches the final reconstruction $g(x)_{0}=\mathcal{I}(x)$ at $t=0$, using steps of size $\delta$.
+This is achieved by inverting the forward process:
+$g(x)_{t}=\mathcal{I}(x)(1-t)+\mathcal{Z}t,\qquad t\in[0,1],$ (3)
+where $\mathcal{Z}$ is sampled from a known distribution matching the shape of $\mathcal{I}(x)$.
+At each step, I-INR estimates the next state by computing the conditional expectation $\mathbb{E}[g(x)_{0}\mid g(x)_{t}]$, which represents the best possible prediction of the original signal given the current state.
+This estimate is then used to update the reconstruction as follows:
+$\hat{g}(x)_{t-\delta}$ $=\mathbb{E}\left[g(x)_{t-\delta}\mid\hat{g}(x)_{t}\right]$ (4)
+$=\frac{\delta}{t}\,\mathbb{E}\left[g(x)_{0}\mid\hat{g}(x)_{t}\right]+\left(1-\frac{\delta}{t}\right)\hat{g}(x)_{t},$
+by repeatedly applying this update, I-INR smoothly transitions from the initial state to the reconstructed signal, leveraging the conditional expectation at each step to improve accuracy.
+I–INR Training.
+To realize the update rule in Eq.
+(4), we train an implicit neural network $f_{\theta}$ (see Figure.
+2(a)) that maps the intermediate state $\tilde{g}(x)_{t}$, the spatial coordinates $x$, and the time index $t$ directly to the clean target $\mathcal{I}(x)$:
+$\min_{\theta}\;\mathbb{E}_{x,\,t,\,n}\left\|f_{\theta}\!\bigl(\tilde{g}(x)_{t},\,x,\,t\bigr)-\mathcal{I}(x)\right\|_{2}^{2},$ (5)
+$\tilde{g}(x)_{t}=(1-t)\,\mathcal{I}(x)+t\,\mathcal{Z}+\varepsilon tn,$ (6)
+where $\mathcal{Z}\sim p(\mathcal{Z})$ and $n\sim\mathcal{N}(0,1)$.
+The small perturbation $\varepsilon tn$ (delbracioinversion) satisfies the regularity requirement, guaranteeing stable reconstruction during inference.
+The full training procedure appears in Algorithm 1.
+I–INR Reconstruction.
+Reconstruction begins from the latent state $\mathcal{Z}$ and proceeds for $1/\delta$ iterations of Eq.  4, iteratively refining the estimate towards the final signal (Algorithm 2).
+Algorithm 1 I-INRs Model Training of $f_{\theta}$
+[…]
+[…]
+3:repeat
+4:  $x\sim p(x)$
+5:  $t\sim\mathcal{U}(0,1),\;n\sim\mathcal{N}(0,I)$
+6:  $\tilde{g}(x)_{t}\leftarrow(1-t)\,\mathcal{I}(x)+t\,\mathcal{Z}+\varepsilon\,t\,n$
+7:  $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\bigl\|f_{\theta}(\tilde{g}(x)_{t},\,x,\,t)-\mathcal{I}(x)\bigr\|_{2}^{2}$
+8:until converged
 Algorithm 2 I-INRs Model Reconstruction
-$\mathcal{Z}$, $f_{\theta}(\cdot,t)$, $\delta$, $x$
-$\hat{g}(x)_{1}=\mathcal{Z}$
-for $t\in\{1,1-\delta,1-2\delta,\dots,0\}$ do $\triangleright$ $t$ decreases from $1$ to 00 in steps of $\delta$
-     $\hat{g}(x)_{t-\delta}=\frac{\delta}{t}f_{\theta}(\hat{g}(x)_{t},{x},t)+\left(1-\frac{\delta}{t}\right)\hat{g}(x)_{t}$
-end for
-return $\tilde{g}(x)_{0}$
-Figure 4: Visual quality comparison of super-resolution results at 2× scale using various methods.
-The ground truth is compared against baseline INR methods SIREN [39], WIRE [36], Gauss [30], and their iterative counterparts.
-The iterative approaches consistently achieve sharper reconstructions with fewer artifacts, effectively preserving finer details and high-frequency structures.
-Network Architecture: Figure 2(a) illustrates the I-INR architecture that integrates an Auxiliary Network and a Backbone (such as SIREN) to compute the output $I(x)_{t}$.
-The input $x$ is processed independently by both the Auxiliary Network and the Backbone, while the Auxiliary Network also incorporates feedback from the previous step.
-The output of the Auxiliary Network is concatenated with the features extracted by the Backbone, forming a joint representation.
-This concatenated feature then passes through a projection layer, which further refines the combined features.
-The final output is obtained as the dot product of the projection layer and the Backbone.
+[…]
+[…]
+3:for $t\leftarrow 1$ to 0 step $\delta$ do
+4:  $\hat{g}(x)_{t-\delta}\leftarrow\dfrac{\delta}{t}\,f_{\theta}\bigl(\hat{g}(x)_{t},x,t\bigr)+\left(1-\dfrac{\delta}{t}\right)\hat{g}(x)_{t}$
+5:end for
+6:return $\hat{g}(x)_{0}$
+[…]
+Network Architecture.
+Figure 2(a) provides an overview of our plug-and-play I-INR framework, which comprises three main components: a Backbone network, a FeedbackNet module, and a FuseNet module.
+The Backbone extracts initial features from the input; the FeedbackNet incorporates the intermediate state and time conditioning; and the FuseNet merges these feature streams.
+The final output is given by Eq.
+(7).
+This design enables flexible integration with existing INR architectures and supports efficient iterative refinement.
+The information flow between components is formalized as
+$f_{\theta}\bigl(\hat{g}(x)_{t},\,x,\,t\bigr)=\operatorname{FuseNet}\bigl(\operatorname{concat}(\mathbf{f},\mathbf{b})\bigr)\odot\mathbf{b},$ (7)
+where
+$\mathbf{b}:=\operatorname{Backbone}(x),\qquad\mathbf{f}:=\operatorname{FeedbackNet}\bigl(\hat{g}(x)_{t},\,x,\,t\bigr).$
+At each iteration, the current state is processed by the FeedbackNet, and its output is fused with the Backbone features via FuseNet.
+Notably, the Backbone is forward-passed only once throughout the reconstruction, while the lightweight FeedbackNet and FuseNet operate at each iteration, resulting in a substantial reduction in computational cost.
 5 Experiments
@@
 Our evaluation encompasses regression tasks, such as 2D and 3D signal fitting, as well as generalization and robustness tasks, including image SR and image denoising.
-To demonstrate its comparative advantage, we benchmark our method against established baseline approaches, including SIREN [39], Gauss [30], and WIRE [36].
-Implementation Settings: The proposed I-INR framework is implemented in PyTorch [25] and optimized using the Adam optimizer [10].
-For all experiments, the noise parameter $\epsilon$ is empirically set to 0.1, and the inference $steps$ are fixed to 2 unless stated otherwise.
-The initial state $\mathcal{Z}$ is sampled from standard normal Gaussian distribution across all experiments.
-All baseline methods, including SIREN [39], WIRE [36], and Gauss [30], as well as their iterative counterparts, are initialized following the configurations specified in their respective papers.
-In I-INR, both the Auxiliary Network and the Projection layer employ lightweight MLPs to ensure computational efficiency while preserving expressivity.
-Evaluation Metrics: To comprehensively evaluate our proposed method, we conduct extensive experiments using both distortion-based metrics, such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) [45], as well as perceptual metrics like Learned Perceptual Image Patch Similarity (LPIPS) [48].
+To demonstrate its comparative advantage, we benchmark our method against established baseline approaches, including SIREN, Gauss, and WIRE.
+Implementation Settings.
+The proposed I-INR framework is implemented in PyTorch (paszke2019pytorch) and optimized using the Adam optimizer (diederik2014adam).
+For all experiments, the noise parameter $\epsilon$ is empirically set to 0.1, and the inference stepssteps are fixed to 2 unless stated otherwise.
+The initial state $\mathcal{Z}$ is sampled from a standard normal Gaussian distribution across all experiments.
+All baseline methods, including SIREN, WIRE, and Gauss, as well as their iterative counterparts, are initialized following the configurations specified in their respective papers.
+In our implementation, both FeedbackNet and FuseNet are lightweight MLPs with two layers, each layer having a width of 30 for FeedbackNet and 100 for FuseNet.
+Evaluation Metrics.
+To comprehensively evaluate our proposed method, we conduct extensive experiments using both distortion-based metrics, such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) (wang2004image), as well as perceptual metrics like Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable).
+For a detailed analysis, refer to the supplementary materials.
 5.1 Results
@@
 Figure 5: Visual comparison of denoising results using various methods.
-The ground truth is compared against noisy input, baseline methods SIREN [39], WIRE [36], and Gauss [30], as well as their iterative counterparts.
+The ground truth is compared against noisy input, baseline methods SIREN, WIRE, and Gauss, as well as their iterative counterparts.
 I-INR demonstrates superior artifact reduction and detail preservation compared to their non-iterative counterparts.
-Table 1: Image fitting results on the Kodak dataset comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts.
-The iterative approach outperforms baseline single-shot architectures by achieving higher reconstruction quality and enhanced fidelity.
-SIREN [39] WIRE [36] Gauss [30]
+SIREN WIRE Gauss
 PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
@@
 Ours 37.53 0.961 33.73 0.924 31.93 0.884
-Image Fitting: For image fitting, we employ SIREN [39], WIRE [36], and Gauss [30], along with their iterative variants.
-Each network utilizes a three-layer MLP with 300 neurons per layer.
-In the case of iterative models, this corresponds to the Backbone INR illustrated in Figure 2(a).
-The networks are trained on the full-resolution Kodak dataset (768$\times$512 or 512$\times$768) [20], and the results, averaged across the entire dataset, are presented in Table 1.
-Our findings demonstrate that the proposed iterative models significantly outperform their one-shot counterparts.
-Figure 3 provides a qualitative comparison of all networks and their iterative adaptations on the kodim19 image.
-The error maps of baseline models exhibit homogeneous noise across the image, with more pronounced errors in high-frequency regions.
-In contrast, the iterative approaches yield smoother reconstructions with reduced errors throughout the image, including high-frequency areas.
-This ultimately results in a substantial improvement in fidelity, with PSNR gains of up to 2.7 dB observed for I-SIREN compared to its one-shot equivalent.
-Image Super-resolution: We evaluate our I-INR framework against traditional one-shot INRs on SR tasks using 40 images from the DIV2K dataset [1].
-The model is trained exclusively on 2×2$\times$2 × SR with a two-layer MLP of 256 neurons per layer.
-To assess generalization, we evaluate I-INR on both 2×2$\times$2 × and 4×4$\times$4 × SR, despite training only on 2×2$\times$2 ×.
-Table 2 compares its performance with single-shot baselines, demonstrating that our approach consistently enhances fidelity (measured as PSNR) and perceptual quality (measured as LPIPS) across all architectures.
-Figure 4 visualizes the butterfly image for 2×2$\times$2 × SR (please refer to supplementary material for visualization at 4×4$\times$4 × scale) , comparing each baseline with its iterative counterpart.
+Table 1: Image fitting results on Kodak comparing SIREN, WIRE, and Gauss with their iterative counterparts.
+Image Fitting.
+For image fitting, each baseline employs a 3-layer MLP with 300 neurons per layer, serving as the Backbone INR for iterative models (Figure 2(a)).
+All networks are trained on the full-resolution Kodak dataset (kodak), with average results reported in Table 1.
+The proposed iterative models consistently outperform their one-shot counterparts.
+As shown in Figure 3, error maps from baseline models display uniform noise, especially in high-frequency regions, whereas iterative variants produce smoother reconstructions with reduced errors, including in challenging areas.
+Image Super-resolution.
+We evaluate our I-INR framework against traditional one-shot INRs on SR tasks using 40 images from the DIV2K dataset (agustsson2017ntire).
+The model is trained exclusively on $2\times$ SR with a two-layer MLP of 256 neurons per layer.
+To assess generalization, we evaluate I-INR on both $2\times$ and $4\times$ SR, despite training only on $2\times$.
+Table 2 compares its performance with single-shot baselines, demonstrating that our approach consistently enhances fidelity (PSNR) and perceptual quality (LPIPS) across all architectures.
+Figure 4 visualizes the butterfly image for $2\times$ SR (refer to supplementary for visualization at $4\times$), comparing each baseline with its iterative counterpart.
 The proposed approach consistently outperforms baseline methods, generating sharper reconstructions with fewer artifacts across all non-linearities.
-This improvement stems from the iterative process, which progressively refines the reconstruction, enabling more precise recovery of high-frequency details and allowing the model to generalize effectively beyond its training scale.
-Table 2: Super-resolution (SR) results for 2×2$\times$2 × and 4×4$\times$4 × evaluations using a model trained exclusively on 2×2$\times$2 × SR.
-The table compares SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts.
-PSNR (higher is better) and LPIPS (lower is better) are reported, with the best result in each column highlighted in bold.
-Scale Method SIREN [39] WIRE [36] Gauss [30]
-Factor PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
+Scale Method SIREN WIRE Gauss
+PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
 $2\times$ Baseline 26.77 0.414 26.14 0.457 25.19 0.538
 Ours 27.64 0.367 27.21 0.388 26.82 0.363
-4$\times$ Baseline 25.03 0.597 24.57 0.6179 23.85 0.673
+$4\times$ Baseline 25.03 0.597 24.57 0.618 23.85 0.673
 Ours 25.53 0.575 25.78 0.496 25.18 0.620
-Table 3: Image denoising results on 40 images sampled from the DIV2K dataset [1], comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts.
-The best values in each column are highlighted in bold.
-SIREN [39] WIRE [36] Gauss [30]
+Table 2: Comparison of (SR) performance for $2\times$ and $4\times$ upscaling using a model trained on $2\times$ SR.
+Results are shown alongside SIREN, WIRE, and Gauss, including their iterative counterparts.
+The best results are shown in bold.
+SIREN WIRE Gauss
 PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
 Ours 25.59 0.540 24.76 0.490 24.20 0.533
+Table 3: Image denoising results on 40 images sampled from the DIV2K dataset, comparing SIREN, WIRE, and Gauss with their iterative counterparts.
 Figure 6: Visualization of 3D occupancy reconstruction results using various methods and their iterative counterparts.
-The ground truth is compared to baseline models (SIREN [39], WIRE [36], Gauss [30]) and their iterative counterparts.
-Iterative methods demonstrate superior reconstruction quality, achieving higher Intersection over Union (IOU) scores and capturing finer geometric details.
-Image Denoising: To assess the robustness of I-INRs in modeling noisy signals, we utilize 40 high-resolution color images from the DIV2K dataset [1] and introduce Poisson-distributed noise to each pixel, with a maximum mean photon count of 30 and a readout noise level of 2, following  [36].
+The ground truth is compared to baseline models SIREN, WIRE, Gauss, and their iterative counterparts.
+Image Denoising.
+To assess the robustness of I-INRs in modeling noisy signals, we utilize 40 high-resolution color images from the DIV2K dataset and introduce Poisson-distributed noise to each pixel, with a maximum mean photon count of 30 and a readout noise level of 2, following  (saragadam2023wire).
 The same architecture and training procedure as SR is employed.
-Table 3 presents quantitative results comparing the denoising performance of SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts.
-Across all three methods, the iterative models consistently achieve significant improvements in both PSNR and LPIPS over their single-shot baselines.
-Figure 5 provides a qualitative comparison of denoising performance for SIREN [39], WIRE [36], and Gauss [30] alongside their iterative versions.
+Table 3 presents quantitative results comparing the denoising performance of baseline architectures with their iterative counterparts.
+Iterative models consistently achieve significant improvements in both PSNR and LPIPS over their single-shot baselines.
+Figure 5 provides a qualitative comparison of denoising performance for SIREN, WIRE, and Gauss alongside their iterative versions.
 The iterative approach significantly enhances reconstruction quality across all architectures, achieving up to +3.25 dB PSNR improvement (for SIREN).
-Refer to caption Refer to caption
-(a) (b)
-Figure 7: On the Kodak dataset for image fitting: (a) PSNR vs. training FLOPs comparison between SIREN [39] and I-SIREN. (b) PSNR vs. training iterations for I-SIREN with different learning rates.
-Figure 8: Analysis of I-SIREN’s performance for super-resolution at 2$\times$ resolution across different iteration $steps$.
-(a) I-SIREN achieves peak PSNR at $steps=4$, where further increasing time steps enhances perceptual quality but reduces fidelity due to the fidelity-perception tradeoff.
-(b) Increasing the number of steps consistently improves perceptual quality.
-(c) PSNR reaches its highest performance at $steps=4$, and while additional iterations slightly reduce PSNR, it remains superior to the baseline.
-3D Occupancy: The 3D occupancy task aims to reconstruct 3D shapes by modeling the occupancy of points in space.
-We conduct experiments on the Armadillo, Dragon, and Thai Statue 3D scenes [12], utilizing a two-layer MLP with 256 neurons per layer for both the baseline INRs and their iterative counterparts.
+[…]
+3D Occupancy.
+The 3D occupancy task aims to reconstruct 3D shapes by modeling the occupancy of points in space.
+We conduct experiments on the Armadillo, Dragon, and Thai Statue 3D scenes (gal2006salient), utilizing a two-layer MLP with 256 neurons per layer for both the baseline INRs and their iterative counterparts.
 Table 4 presents the averaged Intersection over Union (IoU) results, while Figure 6 visualizes the reconstructions achieved using different nonlinearities and their iterative variants.
 Notably, I-INRs outperform their single-shot baselines, generating sharper and more precise reconstructions.
-The iterative approach enhances detail preservation, particularly in complex regions.
-For instance, I-SIREN reconstructs significantly finer details at the base of the statue compared to the baseline SIREN, demonstrating its ability to model intricate spatial structures more effectively.
-Table 4: Averaged Intersection over Union (IoU) results comparing SIREN [39], WIRE [36], and Gauss [30] with their iterative counterparts on the Armadillo, Dragon, and Thai Statue 3D scenes.
-SIREN[39$\uparrow$ WIRE[36$\uparrow$ Gauss[30$\uparrow$
+The iterative approach enhances detail preservation, particularly in complex regions as shown in Figure 6.
+SIREN $\uparrow$ WIRE $\uparrow$ Gauss $\uparrow$
 Baseline 0.9840 0.9917 0.9855
 Ours 0.9934 0.9950 0.9967
+Table 4: Averaged Intersection over Union (IoU) results comparing SIREN, WIRE, and Gauss with their iterative counterparts on the Armadillo, Dragon, and Thai Statue 3D.
 5.2 Ablation Studies
-Initial State: I-INR reconstructs a signal iteratively over multiple steps, starting from an initial state $\mathcal{Z}$.
+We ablate inference complexity, Feedback/FuseNet impact, and reconstruction steps.
+Further analysis on initial state, training cost, layer depth, and statistical significance is included in the supplementary material.
+Inference Complexity.
+During reconstruction, I-SIREN executes the Backbone network only once, while subsequent refinement steps involve forward passes through the lightweight FeedbackNet and FuseNet modules.
+The iterative setup introduces minimal computational overhead: the Backbone accounts for 106.8 GFLOPs, and each refinement step adds just 0.43 GFLOPs.
+For two refinement steps ($\text{steps}=2$), the total computational cost amounts to 107.6 GFLOPs, approximately equivalent to the original SIREN.
+Impact of FeedbackNet and FuseNet:
+To evaluate the impact of architectural components of I-INR, we performed ablation studies by selectively disabling the FeedbackNet and FuseNet modules and varying their parameter scales.
+These experiments were conducted on the image fitting task using I-SIREN on the Kodak dataset.
+As summarized in Table 5, removing both modules leads to a significant performance drop, with PSNR falling to 20.27.
+Introducing FeedbackNet alone substantially improves reconstruction quality, and the addition of FuseNet provides further gains.
+The best performance is achieved when both modules are included, with moderate parameter scaling yielding an optimal balance between computational cost and accuracy.
+FeedbackNet (Params) FuseNet (Params) PSNR
+20.27
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 5: Impact of FeedbackNet and FuseNet on image fitting task with I-SIREN on the Kodak dataset.
+Analysis of Reconstruction Steps.
+We evaluate the effect of reconstruction steps on $2{\times}$ super-resolution using 40 images from the DIV2K dataset (Figures 7b and 7c).
+PSNR peaks at $steps=4$, with the largest improvement from 1 to 2 steps; beyond 2, gains saturate.
+While PSNR slightly drops for $steps>4$, perceptual quality (LPIPS) continues to improve, reflecting the Perception-Distortion Tradeoff (blau2018perception).
+Figure 7a visually compares SIREN and I-SIREN across steps, showing that I-SIREN produces increasingly vivid and detailed reconstructions, confirming that additional refinement steps are beneficial for capturing finer textures and high-frequency components.
+This trend generalizes to other tasks, as discussed in the supplementary material.
+6 Conclusion
+This work introduces a novel I-INR framework that progressively reconstructs a signal.
+The proposed method serves as a plug-and-play solution compatible with existing INR methods, significantly enhancing their performance.
+Through extensive experiments, we demonstrate that I-INR effectively captures high-frequency details, improves noise robustness, and achieves superior reconstruction quality across various tasks and nonlinearities.
+7 Additional Results
+7.1 Additional Quantative Results
+Table 6 presents a detailed comparison of our proposed method against the baseline across image fitting, denoising, and SR at $2\times$ for all three evaluation metrics (PSNR, SSIM, and LPIPS).
+Our iterative approach consistently enhances performance across tasks and different baselines, achieving superior reconstruction quality.
+7.2 Additional Qualitative Comparison
+We also present an additional qualitative comparison of our proposed method against SIREN, Gauss, and WIRE.
+Figures 891011, illustrates a comparative analysis for image fitting, SR at $2\times$ and $4\times$ scales, and image denoising, demonstrating the improvements achieved by our approach over baseline INRs.
+Our proposed method consistently outperforms single-shot baselines, demonstrating superior reconstruction quality across various models and tasks.
+8 Additional Ablation Studies
+Initial State.
+I-INR reconstructs a signal iteratively over multiple steps, starting from an initial state $\mathcal{Z}$.
 The choice of this initial state can significantly impact the reconstruction process.
@@
 The results, summarized in Table 7, highlight the superiority of noise-based initial states in achieving higher PSNR and SSIM values.
-Table 5: Impact of different initial states ($\mathcal{Z}$) on image fitting performance using the I-INR framework on the Kodak dataset.
-Noise-based initial states consistently achieve the best results across various non-linearities.
-I-SIREN I-WIRE I-Gauss
-PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-Noise 37.53 0.961 33.73 0.924 31.93 0.884
-Ones 37.42 0.961 33.33 0.921 30.96 0.867
-Zeros 37.47 0.961 33.39 0.921 31.22 0.882
-Training and Inference Complexity: Our proposed I-INR method achieves faster convergence than the non-iterative baseline while maintaining similar training complexity, as shown in Figure 7a.
+Training Complexity.
+Our proposed I-INR method achieves faster convergence than the non-iterative baseline while maintaining similar training complexity, as shown in Figure 12a.
 The results are averaged over the Kodak dataset for the image fitting task.
 Figure 12b further illustrates I-SIREN’s training performance under different learning rates.
-While I-SIREN converges faster, this improvement comes at a slightly higher per-iteration computational cost, requiring 107.2 GFLOPS compared to 106.8 GFLOPS for SIREN.
-Additionally, since I-SIREN reconstructs images iteratively, the total inference cost scales with the number of inference steps.
-In our experiments, we consistently use $steps=2$, resulting in a total inference complexity of 214.4 GFLOPS for I-SIREN, compared to 106.8 GFLOPS for SIREN.
-Number of Layers: For the image fitting task on the Kodak dataset, we investigate the impact of increasing the number of layers in SIREN and its iterative counterpart, I-SIREN.
+Number of Layers in Backbone Network.
+For the image fitting task on the Kodak dataset, we investigate the impact of increasing the number of layers in SIREN and its iterative counterpart, I-SIREN.
 The results, presented in Table 8, show that while the performance of SIREN saturates beyond a certain depth, specifically at five layers, I-SIREN continues to improve as more layers are added.
 Notably, a three-layer I-SIREN outperforms a five-layer SIREN, demonstrating the efficiency of the iterative approach.
-Table 6: Impact of varying layer depths on the image fitting task using SIREN [39] and I-SIREN on the Kodak dataset.
-I-SIREN consistently outperforms SIREN [39], with a three-layer I-SIREN achieving higher PSNR than a five-layer SIREN [39].
-Layers I-SIREN SIREN
-PSNR Params (k) PSNR Params (k)
-3 37.53 274 34.62 273
-4 38.71 364 36.60 363
-5 39.01 455 36.49 453
-Analysis of Reconstruction Steps: We analyze the impact of reconstruction steps on 2$\times$ super-resolution using 40 images sampled from the DIV2K dataset, as shown in Figures 8b and 8c.
-Our findings reveal that the model achieves its highest PSNR at $steps=4$, with the most significant gain occurring when $steps$ increases from 1 to 2.
-Beyond $steps=2$, the PSNR begins to plateau.
-Notably, although PSNR decreases for $steps>4$, the perceptual quality (measured by LPIPS) continues to improve.
-This behavior aligns with the Perception-Distortion Tradeoff [2], where introducing additional high-frequency details enhances realism but slightly deviates from the ground-truth, thus lowering distortion-based metrics such as PSNR.
-Figure 8a presents a visual comparison of SIREN and I-SIREN on an image from DIV2K dataset at various time steps.
-I-SIREN reconstructions become progressively more vibrant and perceptually appealing as step increases, confirming that additional refinement steps are beneficial for capturing finer textures and high-frequency components.
-This trend holds for other reconstruction tasks as well, as detailed in our supplementary material.
-5 Conclusion
-This work introduces a novel I-INR framework that progressively reconstructs a signal.
-The proposed method serves as a plug-and-play solution compatible with existing INR methods, significantly enhancing their performance.
-Through extensive experiments, we demonstrate that I-INR effectively captures high-frequency details, improves noise robustness, and achieves superior reconstruction quality across various tasks and nonlinearities.
-Supplementary Material
-6 Additional Results
-6.1 Additional Quantative Results
-Table 7 presents a detailed comparison of our proposed method against the baseline across image fitting, denoising, and SR at 2×2$\times$2 × for all three evaluation metrics (PSNR, SSIM, and LPIPS).
-Our iterative approach consistently enhances performance across tasks and different baselines, achieving superior reconstruction quality.
-6.2 Additional Qualitative Comparison
-We also present an additional qualitative comparison of our proposed method against SIREN, Gauss, and WIRE.
-Figures 9101112, illustrates a comparative analysis for image fitting, SR at 2×2$\times$2 × and 4×4$\times$4 × scales, and image denoising, demonstrating the improvements achieved by our approach over baseline INRs.
-Our proposed method consistently outperforms single-shot baselines, demonstrating superior reconstruction quality across various models and tasks.
-7 Additional Ablation Studies
-Additional Reconstruction Steps Analysis: For SR at 2×2$\times$2 ×, we further analyze the reconstruction quality of I-SIREN over multiple steps, comparing it to SIREN and the ground truth.
+Additional Reconstruction Steps Analysis.
+For SR at $2\times$, we further analyze the reconstruction quality of I-SIREN over multiple steps, comparing it to SIREN and the ground truth.
 As shown in Figure 13, the results demonstrate that with an increasing number of steps, details become more pronounced, and more high-frequency information is incorporated into the reconstruction.
@@
 We also analyze the effect of reconstruction steps on image denoising performance using 40 images from the DIV2K dataset (see Figure 15).
-Our findings indicate that PSNR peaks at $steps=2$ before gradually declining as $steps$ increases.
+Our findings indicate that PSNR peaks at $steps=2$ before gradually declining as stepssteps increases.
 Meanwhile, LPIPS improves until $step=4$ and stabilizes around step 5.
@@
 At higher iterations, the model inadvertently reconstructs these distortions, leading to a decline in performance.
-Feature Fusion: We conduct ablation studies on the final feature-fusion mechanism for signal reconstruction, evaluating two strategies that combine the outputs of the Backbone and Projection modules (Figure 2a of main manuscript):
-[…]
-[…]
-In Multiplicative fusion, the final output $\mathcal{I}(x)_{t}$ is obtained via element-wise multiplication of the Backbone and Projection outputs.
-The Adaptive method, conversely, uses the interpolation factor $t$ to smoothly shift emphasis from the Backbone output (dominant at $t=1$) toward the Projection output at higher values of $t$.
+Feature Fusion.
+We conduct ablation studies on the final feature-fusion mechanism for signal reconstruction, evaluating two strategies that combine the outputs of the Backbone and FuseNet modules (Figure 2a of main manuscript):
+$\mathbf{b}:=\operatorname{Backbone}(x),\qquad\mathbf{f}:=\operatorname{FeedbackNet}(\hat{g}(x)_{t},\,x,\,t)$
+$\mathbf{z}:=\operatorname{FuseNet}(\operatorname{concat}(\mathbf{f},\,\mathbf{b}))$
+$\text{Multiplicative:}\quad\mathit{out}=\mathbf{b}\odot\mathbf{z},$
+$\text{Adaptive:}\quad\mathit{out}=\mathbf{b}\cdot t+\mathbf{z}\cdot(1-t)$
+In Multiplicative fusion, the final output is obtained via element-wise multiplication of the Backbone and FuseNet outputs.
+The Adaptive method, conversely, uses the interpolation factor $t$ to smoothly shift emphasis from the Backbone output (dominant at $t=1$) toward the FuseNet output at higher values of $t$.
 Quantitative results comparing both fusion strategies for the image-fitting task are summarized in Table 9.
@@
 Therefore, we adopt Multiplicative fusion as the standard approach for our method.
-Table 7: Experimental results for Denoising, Image Fitting, and Super Resolution for Baseline model with Ours, evaluating their performance across three tasks.
-The best values in each column are highlighted in bold.
+Time Complexity.
+We compare SIREN and I-SIREN on the Kodak dataset, trained for 2000 iterations on an NVIDIA RTX 4090 GPU.
+As shown in Table 10, I-SIREN adds only a small overhead in training time ($6\%$) and inference latency (+3+3 ms per image) while achieving higher reconstruction quality, confirming its efficiency.
+Statistical Significance.
+All prior experiments were conducted using a single random seed.
+To assess the robustness and statistical significance of our approach, we additionally evaluate performance across five different seeds.
+Table 11 reports the results for image fitting, SR, and denoising tasks, comparing SIREN and I-SIREN, along with confidence intervals.
+The dataset and experimental settings remain consistent with those previously described.
+As shown in Table 11, I-SIREN consistently outperforms SIREN across all tasks.
+The improvements achieved by I-SIREN are statistically significant, with p-values << 0.00001 based on the Wilcoxon signed-rank test.
 SIREN Gauss WIRE
@@
 Ours 25.59 0.619 0.540 24.20 0.612 0.533 24.76 0.645 0.490
-Table 8: Comparison of fusion methods for the image-fitting task at $steps=2$ on the Kodak dataset with Gaussian initial state ($\mathcal{Z}\sim\mathcal{N}(0,1)$).
+Table 6: Experimental results for Denoising, Image Fitting, and Super Resolution for Baseline model with Ours, evaluating their performance across three tasks.
+The best values in each column are highlighted in bold.
+I-SIREN I-WIRE I-Gauss
+PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+Noise 37.53 0.961 33.73 0.924 31.93 0.884
+Ones 37.42 0.961 33.33 0.921 30.96 0.867
+Zeros 37.47 0.961 33.39 0.921 31.22 0.882
+Table 7: Impact of different initial states ($\mathcal{Z}$) on image fitting task using the I-INR framework on Kodak dataset.
+Layers I-SIREN SIREN
+PSNR Params (k) PSNR Params (k)
+3 37.53 274 34.62 273
+4 38.71 364 36.60 363
+5 39.01 455 36.49 453
+Table 8: Impact of varying Backbone network layer depths on the image fitting task using SIREN and I-SIREN on the Kodak dataset.
 I-SIREN I-WIRE I-Gauss
@@
 Adaptive 37.01 0.960 0.037 33.63 0.923 0.078 31.80 0.883 0.129
+Table 9: Comparison of fusion methods for the image-fitting task at $steps=2$ on the Kodak dataset with Gaussian initial state ($\mathcal{Z}\sim\mathcal{N}(0,1)$).
+Method Train Time (s) Inference (ms / image) PSNR (dB)
+SIREN 100 18.9 34.57
+I-SIREN (2 steps) 106 21.9 37.53
+Table 10: Training and inference time comparison on the Kodak dataset (RTX 4090).
+Method Image Fitting SR Denoising
+I-SIREN 37.50 $\pm$ 0.247 27.22 $\pm$ 0.045 25.36 $\pm$ 0.044
+SIREN 34.59 $\pm$ 0.027 26.77 $\pm$ 0.035 23.89 $\pm$ 0.021
+Table 11: Evaluation on 5 different seeds for image fitting, super-resolution tasks, and denoising on SIREN and I-SIREN.
 9 Training Setup
@@
 Additionally, we used 2000 training iterations for image fitting, super-resolution (SR), and denoising tasks, while for object occupancy, we used 200 training iterations.
+Task Parameter I-Gauss I-SIREN I-WIRE
+Image Fitting $\omega$ - 52 7
+$\sigma$ 18 - 13
+Image SR $\omega$ - 30 4
+$\sigma$ 11 - 10
+Image Denoising $\omega$ - 55 10
+$\sigma$ 18 - 16
+3D Occupancy $\omega$ - 55 10
+$\sigma$ 17 - 20
 Table 12: Hyperparameter settings for I-INRs across different tasks.
-Task Parameter I-Gauss I-SIREN I-WIRE
-Image Fitting $\omega$ 52 7
-$\sigma$ 18 - 13
-Image SR $\omega$ 30 4
-$\sigma$ 11 - 10
-Image Denoising $\omega$ 55 10
-$\sigma$ 18 - 16
-3D Occupancy $\omega$ 55 10
-$\sigma$ 17 - 20
 Figure 8: Image fitting results for different non-linearities and their iterative extensions (prefix ”I”).
-Figure 10: Visual quality comparison of super-resolution results at 2$\times$ scale using various methods.
-The ground truth is compared against baseline INR methods SIREN, WIRE, Gauss, and their iterative counterparts.
-The iterative approaches consistently achieve sharper reconstructions with fewer artifacts, effectively preserving finer details and high-frequency structures.
-Figure 11: Visual quality comparison of super-resolution results at 4$\times$ scale using various methods.
-The ground truth is compared against baseline INR methods SIREN, WIRE, Gauss, and their iterative counterparts.
-The iterative approaches consistently achieve sharper reconstructions with fewer artifacts, effectively preserving finer details and high-frequency structures.
+[…]
+[…]
 Figure 11: Visual comparison of denoising results using various methods.
@@
 I-INR demonstrates superior artifact reduction and detail preservation compared to their non-iterative counterparts.
-Figure 13: Analysis of I-SIREN’s performance for super-resolution at 2$\times$ resolution across different steps.
-I-SIREN achieves peak PSNR at $steps=4$, where further increasing steps enhances perceptual quality.
+Refer to caption Refer to caption
+(a) (b)
+Figure 12: On the Kodak dataset for image fitting: (a) PSNR vs. training FLOPs comparison between SIREN and I-SIREN. (b) PSNR vs. training iterations for I-SIREN with different learning rates.
+[…]
 Refer to caption Refer to caption